Obtain the inverse Laplace transform of t cos pt, p is not equal to 0.


1
Expert's answer
2021-01-05T14:06:32-0500

"F(t)=tcos(pt),p\neq0\\
\therefore \mathcal{L}^{-1}\{F(s)\}\\
=\frac{d}{dt}\left(\mathcal{L}^{-1}\left\{\cos \left(ps\right)\right\}\right)+\mathcal{L}^{-1}\left\{\cos \left(ps\right)\right\}\left(0\right)"

Now "\mathcal{L}^{-1}\left\{\cos \left(ps\right)\right\} : No \; solution \;for\; standard \; function"


"\mathcal{L}^{-1}\{F(s)\}" has no solution.


Now for Laplace transform,

"F(t)\\
=tcos(pt)\\
=\frac{1}{2p}2pcos(pt)\\
=\frac{1}{2p}\left[(sin(pt)+pcos(pt))-(sin(pt)-ptcos(pt))\right]"

"\\\;\\
\therefore \mathcal{L}\{F(t)\}\\
=\mathcal{L}\left(\frac{1}{2p}\left[(sin(pt)+pcos(pt))-(sin(pt)-ptcos(pt))\right]\right)\\
=\frac{1}{2p}\left[\mathcal{L}(sin(pt)+pcos(pt))-\mathcal{L}(sin(pt)-ptcos(pt))\right]\\
= \frac{1}{2p}\left[\frac{{2p{s^2}}}{{{{\left( {{s^2} + {p^2}} \right)}^2}}}-\frac{{2{p^3}}}{{{{\left( {{s^2} + {p^2}} \right)}^2}}}\right]\\
=\frac{1}{2p}\times \frac{{2ps^2-2{p^3}}}{{{{\left( {{s^2} + {p^2}} \right)}^2}}}
=\frac{1}{2p}\times \frac{{2p(s^2-p^2)}}{{{{\left( {{s^2} + {p^2}} \right)}^2}}}\\
=\frac{s^2-p^2}{(s^2+p^2)^2}"


Therefore, Laplace transform of "\pmb{F(t)=\frac{s^2-p^2}{(s^2+p^2)^2}}" .